# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v1

@@ Abstract.
We formalized a complete proof of the Auslander–Buchsbaum–Serre criterion in the Lean4 theorem prover. For a local ring, rather than following the well-known proof that considers the residue field as the quotient of the ring by a regular sequence to compute projective dimension and uses the Koszul complex to show the dimension of the cotangent space is at most the global dimension, we prove the criterion via maximal Cohen–Macaulay modules and a weakened version of the Ferrand–Vasconcelos theorem, which is more amenable to the formalization process and the current development of mathlib. Our formalization includes the construction of depth and of Cohen–Macaulay modules and rings, which are used frequently in the proof of the criterion. We also developed related results, including the unmixedness theorem for Cohen–Macaulay rings and Hilbert’s Syzygy theorem.
@@ 1. Introduction
The Auslander–Buchsbaum–Serre criterion is a beautiful piece of mathematics which states that for a Noetherian local ring $R$, $R$ is a regular local ring if and only if it has finite global dimension. This criterion did not appear all at once: initially only the “only if” direction was known. At that time, regularity of the localization of a regular local ring at a prime ideal was still an open problem. It was not until 1955 that Jean–Pierre Serre gave the proof of the converse, completing the homological characterization of regular local rings. With this characterization in hand, it is easy to prove that the localization of a regular local ring is again regular. This is one of the earliest applications of homological methods to prove a non-homological result. Furthermore, as a corollary, regularity at all maximal ideals and at all prime ideals coincides, giving a clearer understanding of regular rings.
This project began with the goal of developing the theories of depth and Cohen–Macaulay rings. As the concepts of regular element and regular sequence were already formalized, this goal was within reach given the current library mathlib [7]. The effort was not only of mathematical interest, but also a test of the newly developed homological algebra in mathlib via its application to commutative algebra. As our focus shifted toward combination with homological algebra, we set Hilbert’s Syzygy theorem and the Auslander–Buchsbaum–Serre criterion as targets. However, the $\operatorname{Tor}$ functor was still largely missing from mathlib, and development of the Koszul complex was also incomplete, so the most well-known proof failed on both fronts. The turning point was finding an approach to the global dimension of a regular local ring by proving that a maximal Cohen–Macaulay module over a regular local ring is free; then the target became attainable. Finally, by observing a proof of the Ferrand–Vasconcelos theorem (see [1, Theorem 2.2.8]) showing that induction on dimension can still be performed when focusing only on the unique maximal ideal, we completed the full Auslander–Buchsbaum–Serre criterion. The value of this work is not only in the final theorems: throughout the process we enriched basic constructions in commutative algebra and homological algebra and tested their design and lemmas directly in subsequent developments, making them more user-friendly. We also plan to upstream most of our work into mathlib in the future.
@@
In Section  2 we introduce preliminary results needed for the formalization, including the existing theories in mathlib and the supplementary results we added. We then develop results about depth and Cohen–Macaulay modules and rings in Sections  3 and  4, following [4, Section 6]. This was the earliest part to be formalized. For the theory of regular local rings, Section  5 gives the definition of a regular local ring and basic results about it, such as that a regular local ring is a domain. The Auslander–Buchsbaum–Serre criterion is split into two subsections:  5.2 focuses on proving that the global dimension of a regular local ring equals its Krull dimension, and  5.3 proves a weakened version of the Ferrand–Vasconcelos theorem that implies the converse. These two subsections roughly follow a modified proof obtained from combining [6] and [1, Section 2.2]. In subsection  5.4 we state our final results and their corollaries. In Section  6 we describe implementation issues, related work, and directions for future work.
@@
Throughout the paper, we use the symbol W for external links. Most statements and definitions are accompanied by such a link directly to the source code for the corresponding statement in mathlib or in a branch of our fork W. To keep the links stable, the links to the latter point to a fixed commit of the master branch (shortly after cleanup of the project); other links point to the documentation of mathlib. In some code excerpts the dot notation is exhibited: for example, for I of type Ideal R, the call Ideal.depth I M can be shortened to I.depth M.
@@ Acknowledgements
We would like to thank Professor Liang Xiao for the preliminary review and for organizing the grouped work that made this project possible. We also extend our gratitude to Dr. HuanHuan Yu for producing the blueprint of the formalization project. We are also deeply grateful to all the contributors and reviewers who have been involved in the development and update process of both Lean and mathlib, whose efforts have made our work possible. Special thanks go to Professor Joël Riou for developing the derived category and the $\operatorname{Ext}$ functor, which enabled the homological methods. This work is supported in part by National Key R&D Program of China grant 2024YFA1014000.
@@ 2. Preliminaries
Introduction:
In this section, we review some preliminary results that will be used in later proofs.
@@ 2.1. Associated Primes
At the start of this project, we only had the definition associatedPrimes W and basic facts about associated primes in mathlibW. Thanks to the work of Jinzhao Pan, we have the finiteness of associated primes for a finitely generated module over a Noetherian ring in mathlibW.
@@
Combined with the fact that the complement of the non-zero-divisors of a module is exactly the union of its associated primesW, this lemma lets us apply prime-avoidance arguments when looking for a regular element in an ideal.
We also formalized results on associated primes of localized modulesW. For $S$ a multiplicative set of a Noetherian ring $R$ and $M$ a finitely generated $R$-module, one has $\mathrm{Ass}(S^{-1}M)=\mathrm{Ass}(M)\cap\operatorname{Spec}(S^{-1}R)$. Denoting $R^{\prime}$ and $M^{\prime}$ the localizations of $R$ and $M$ at $S$ respectively, the main statement is:
@@
Notice that the induced map $\operatorname{Spec}(R^{\prime})\to\operatorname{Spec}(R)$ is injective, this lemma already states the strongest result.
Furthermore, applying the above to the localization at $p$ a minimal prime over $\mathrm{Ann}(M)$: since associated primes of $M_{p}$ exist, any such associated prime must be the maximal ideal of $R_{p}$. From the previous result we deduce that minimal primes over $\mathrm{Ann}(M)$ lies in $\mathrm{Ass}(M)$W.
@@ 2.2. Krull Dimension of Module
We define the Krull dimension of a module simply as the Krull dimension of its support viewed as a preorder setW. Some basic results are formalized hereW.
Similar to dimension of ring, for $R$-module $M$ quotienting by an element in the jacobson radical of $\mathrm{Ann}(M)$ reduces the dimension by at most oneW.
theorem supportDim_le_supportDim_quotSMulTop_succ_of_mem_jacobson {x : R} (h : x (annihilator R M).jacobson) :
supportDim R M supportDim R (QuotSMulTop x M) + 1
The formalization of this lemma follows [3, Proposition 2.3.4] via [3, Lemma 2.3.3]W.
If we quotient by an element $x$ that is not contained in any minimal prime over $\mathrm{Ann}(M)$, the dimension decreases by at least oneW:
theorem supportDim_quotSMulTop_succ_le_of_notMem_minimalPrimes {x : R} (hn : p (annihilator R M).minimalPrimes, x p) :
supportDim R (QuotSMulTop x M) + 1 supportDim R M
Combining this with the fact that an $M$-regular element is not contained in any minimal prime over $\mathrm{Ann}(M)$, we obtainW:
@@
If $R$ is local, induction on the length of a regular sequence shows that quotienting by a regular sequence decreases the dimension of a module by exactly its lengthW:
@@ 2.3. Dimension of Polynomial Rings
As we need to work with polynomials over Cohen–Macaulay and regular rings, some results about the heights of ideals in polynomial rings over Noetherian rings are required. Let $R$ be a Noetherian ring, we establish the following results, which follow easily from Krull’s height theoremW.
lemma height_map_maximalIdeal [IsLocalRing R] : ((maximalIdeal R).map C).height = (maximalIdeal R).height
lemma height_of_comap_eq_maximalIdeal [IsLocalRing R] (m : Ideal R[X]) [h : m.IsPrime]
(eqm : m.comap C = maximalIdeal R) : m.height (maximalIdeal R).height + 1
lemma height_le_height_comap_succ (p : Ideal R[X]) [p.IsPrime] : p.height (p.comap C).height + 1
From these we obtain the standard result on the Krull dimension of polynomial ring over a Noetherian ringW.
@@
There is an ongoing effort to formalize similar statements more generally by studying heights under ring homomorphisms with the going-down property.(SeeWW.) Because that work is more general and these lemmas here are proven specifically for polynomial rings (with some proofs “forced through”), they are intended as temporary and will be deprecated once the more general versions are available in mathlib. See  6.2 for details.
@@ 2.4. Regular Sequence Under Flat Base Change
As we need to handle preservation of regular sequences under localization, we formalized the more general results that weakly regular sequences are preserved under flat base changeW and regular sequences are preserved under faithfully flat base changeW.
@@
Specializing to the case where the module is the ring itself yields a more user-friendly versionWW.
Finally, we also added specialized lemmas for localized modules and localization, which show that a regular sequence contained in the prime being localized at remains regular after localizationWW.
@@ 2.5. Projective Dimension
As the $\operatorname{Ext}$ functor is a relatively recent addition to the development of mathlib (see [5]), many related constructions are not yet fully available. We only have a predicate describing that an object $X$ of an abelian category $C$ has projective dimension less than some $n\in\mathbb{N}$, defined as the vanishing of all $\operatorname{Ext}_{i}(X,Y),\forall Y,\forall n\leq i$
@@
With only this definition, some operations are awkward (for example adding a number). Even if we try to pick out the minimal $n$ such that HasProjectiveDimensionLT X (n + 1), the disadvantages are significant: first such number may doesn’t exist at all (if $X$ has no finite projective resolution); second, finding it with Nat.find complicates operations and requires decidable instances. For these reasons we want projective dimension as a concrete numeric object.
However, projective dimension cannot be a plain nat: first, we need to give objects with no finite projective resolution value $\infty$, second, by common convention the projective dimension of the zero object is -$\infty$. We therefore define projective dimension in WithBot ℕ∞:
variable {C : Type u} [Category.{v, u} C] [Abelian C]
noncomputable def projectiveDimension (X : C) : WithBot ℕ∞ :=
@@
Although this definition looks a little verbose, it behaves as in the usual conventions. (Out side code excerpts we denote projectiveDimension by $\operatorname{projdim}$)
With this definition we can prove the equivalence projectievDimension X < n HasProjectiveDimensionLT X n and many other useful lemmas to make the numeric version convenient in practice. Many lemmas related to HasProjectiveDimensionLT now have corresponding statements for projectievDimension.
We also prove a lemma specific to the category of $R$-modulesW.
@@
For convenience in formalization, we do not directly prove that $\operatorname{Ext}$ commutes with localization. Noting that the left-hand side being -$\infty$ is equivalent to $M=0$ and similarly for the right-hand side -$\infty$, we instead prove that for each natural number $n$, the statement left hand side no greater than $n$ is equivalent to the right hand side no greater than $n$, by induction on $n$. Then we construct an exact sequence $0\rightarrow N\rightarrow F\rightarrow M\rightarrow 0$ with $F$ free, localizing at a prime $\mathfrak{p}$ yields an exact sequence $0\rightarrow N_{\mathfrak{p}}\rightarrow F_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}\rightarrow 0$. Using the existing lemma about projective dimension in short exact sequencesW, we reduce the case $n+1$ to the case $n$. We are then left to show that $M$ is projective if and only if each localization at prime $M_{\mathfrak{p}}$ is projective, which is already in mathlib WW. There is also a version of this lemma for running only over all maximal ideals W.
Similarly we prove the following lemmaW.
@@
For $S$ a multiplicative set of $R$ and $R$-module $M$, proective dimension of $S^{-1}M$ over $S^{-1}R$ is no greater than projective dimension of $M$ over $R$.
lemma projectiveDimension_le_projectiveDimension_of_isLocalizedModule [Small.{v} R]
(S : Submonoid R) (M : ModuleCat.{v} R) :
projectiveDimension (M.localizedModule S) projectiveDimension M

### v2

@@ Abstract.
We present a comprehensive formalization in the Lean4 theorem prover of the Auslander–Buchsbaum–Serre criterion, which characterizes regular local rings as those Noetherian local rings with finite global dimension. Rather than following the well-known proof that computes the projective dimension of the residue field via quotient by regular sequences and uses the Koszul complex to bound the cotangent space dimension by the global dimension, our approach is built systematically on the formalization of depth defined via the vanishing of Ext functors. We establish key homological results including Rees’ theorem, the Auslander–Buchsbaum formula, and Ischebeck’s theorem, and further develop the theories of Cohen–Macaulay modules and rings, including a complete formalization of the unmixedness theorem for Cohen–Macaulay rings. To prove the Auslander–Buchsbaum–Serre criterion, we show that maximal Cohen–Macaulay modules over regular local rings are free and establish a weakened form of the Ferrand–Vasconcelos theorem specific for the unique maximal ideal. As corollaries, we deduce that regularity can be checked at maximal ideals and formalize Hilbert’s Syzygy Theorem. This work demonstrates how homological algebra can be effectively employed in the formalization of commutative algebra, providing extensive infrastructure for future developments in the field.
@@ 1. Introduction
The Auslander–Buchsbaum–Serre criterion is a beautiful mathematical result stating that for a Noetherian local ring $R$, the ring $R$ is regular if and only if it has finite global dimension. This criterion was not proved all at once. At first, only the “only if” direction was known, as shown in [3]. At that time, it was still an open problem whether the localization of a regular local ring at a prime ideal remains regular. The homological characterization of regular local rings was completed only after Jean–Pierre Serre proved the converse direction later in [20], as formally recorded in [1]. Once this characterization is available, it becomes straightforward to prove that the localization of a regular local ring is again regular. This is one of the earliest examples where homological methods were used to prove a result that is not itself homological. It shows how homological invariants—such as projective dimension, depth, and homological dimension—can detect and control purely “geometric” or “structural” properties of rings. This opened the door for homological algebra to become a central tool in commutative algebra and algebraic geometry. A similar homological perspective later led to the Auslander–Buchsbaum formula, Serre’s intersection multiplicity theory [21], the development of Cohen–Macaulay and Gorenstein ring theory, and deformation theory, for example [9].
This project began with the goal of developing the theories of depth and Cohen–Macaulay rings. As the concepts of regular element and regular sequence were already formalized, this goal was within reach using the current library mathlib [15]. The effort was not only of mathematical interest, but also a test of the newly developed homological algebra in mathlib via its application to commutative algebra. As we shifted our focus toward combining it with homological algebra, we set Hilbert’s Syzygy theorem and the Auslander–Buchsbaum–Serre criterion as our main targets. However, at the time of this work, the $\operatorname{Tor}$ functor was still largely missing from mathlib, and the development of the Koszul complex was also incomplete, so the most well-known proof could not be used in the formalization. The turning point was finding an approach to the global dimension of a regular local ring by proving that a maximal Cohen–Macaulay module over a regular local ring is free; then the target became attainable. Later, inspired by the proof of the Ferrand–Vasconcelos theorem in [5, Theorem 2.2.8], we realized that induction on span rank (minimal number of generators) could still be carried out by focusing only on the unique maximal ideal. This allowed us to complete the full Auslander–Buchsbaum–Serre criterion. The value of this work lies not only in the final theorems: throughout the process, we formalized more basic constructions in commutative and homological algebra (for example, the definitions of projective dimension and $\operatorname{depth}$), tested these definitions in later developments, and improved their interfaces to make them easier to use. However, some of these additional interfaces are not included in this paper due to space limitations. Currently part of subsection  2.1,  2.5 and all of subsection  2.4,  2.2 are already in mathlib. We plan to upstream most of our work into mathlib in the future.
@@
In Section  2, we introduce the preliminary results needed for the formalization, including existing theories in mathlib and supplementary results we developed during the study of Cohen–Macaulay rings and regular local rings. We then present the development of depth and Cohen–Macaulay modules and rings in Sections  3 and  4, following [16, Section 6]. This was the earliest part to be formalized.
The key preliminary ingredient of the Auslander–Buchsbaum–Serre criterion is the Auslander–Buchsbaum formula (Theorem  3.8) in Subsection  3.2.
Theorem 1.1 (Auslander–Buchsbaum).
Let $(R,\mathfrak{m})$ be a Noetherian local ring and let $M$ be a non-trivial finitely generated $R$-module. If projective dimension of $M$ is finite, then
$\operatorname{projdim}(M)+\operatorname{depth}(M)=\operatorname{depth}(R)$
After developing the basic theory of regular local rings at the beginning of Section  5, and establishing that a regular local ring is a domain in Subsection  5.1, the proof of the criterion splits into two largely independent parts, corresponding to the two directions of the equivalence.
  • In subsection  5.2, we prove that a regular local ring has global dimension equal to its Krull dimension, which establishes one direction of the criterion.
Theorem 1.2.
For regular local ring $R$, $\operatorname{gldim}(R)=\dim(R)$.
Let $(R,\mathfrak{m},k)$ be a regular local ring of dimension $d$. By Theorem  1.1, the projective dimension of any finitely generated module is either infinite or at most $\operatorname{depth}(R)=d$. Since $\operatorname{depth}(k)=0$, if $k$ has finite projective dimension, then again by Theorem  1.1, we have $\operatorname{projdim}(k)=d$. This reduces the problem to showing that all finitely generated modules over a regular local ring have finite projective dimension.
By basic inequalities of $\operatorname{depth}$, we obtain the following:
Definition 1.3.
A module $M$ over local ring $R$ is a maximal Cohen–Macaulay module if $\operatorname{depth}(M)=\dim(R)$.
Proposition 1.4.
The $d$-th syzygy of a finitely generated module $M$ in a free resolution is either a maximal Cohen–Macaulay module or 0
The next result about maximal Cohen–Macaulay modules then ensures that this yields a finite projective (free) resolution of $M$.
Theorem 1.5 ([22, 00NT]).
For regular local ring $R$, any finitely generated maximal Cohen–Macaulay module over $R$ is free.
If an element $x$ is both $R$-regular and $M$-regular, then $M/xM$ is a maximal Cohen–Macaulay module over $R/(x)$ by the following proposition. This allows Theorem  1.5 to be proved by induction on $\dim(R)$.
Proposition 1.6.
For $R$-modules $M$ and $N$ and an $M$-regular element $x\in\mathrm{Ann}(N)$, we have
$\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$.
  • For the converse direction–showing regularity from finite global dimension, we initially attempted to use the following theorem (see Subsection  5.3).
Theorem 1.7 (Ferrand–Vasconcelos, [5, Theorem 2.2.8]).
For an ideal $I$ of a Noetherian local ring $R$, if $I$ has finite projective dimension and $I/I^{2}$ is free over $R/I$, then $I$ is generated by a regular sequence.
However, since some techniques in its proof are not yet fully formalized, we focus only on the case where $I$ is the unique maximal ideal, which still suffices for our goal. Note that a regular sequence in $\mathfrak{m}$ has length at most $\operatorname{depth}(R)\leq\dim(R)$. If $\mathfrak{m}$ has finite projective dimension, the above theorem directly implies that $R$ is regular.
Returning to the case $I=\mathfrak{m}$, the proof roughly follows a modified version combining ideas from [22] and [5, Section 2.2]. By induction on the minimal number of generators of $\mathfrak{m}$, if $x$ is a regular element in $\mathfrak{m}\setminus\mathfrak{m}^{2}$, then $\mathfrak{m}/(x)$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$. The following proposition then implies that $\mathfrak{m}/(x)$ has finite projective dimension, so the induction hypothesis applied to $R/(x)$ completes the proof.
Proposition 1.8.
For $R$-module $M$ and an element $x$ being both $R$-regular and $M$-regular, then
$\operatorname{projdim}_{R/(x)}(M/xM)=\operatorname{projdim}_{R}(M)$
It is worth noting that many properties of Cohen–Macaulay rings are established independently, and not all results presented here are directly tied to the Auslander–Buchsbaum–Serre criterion. However, most of these unrelated results turn out to be useful in proving the unmixedness theorem for Cohen–Macaulay rings.
Theorem 1.9.
A Noetherian ring $R$ is Cohen–Macaulay if and only if for any ideal $I=(r_{1},\cdots,r_{k})$ with $0pt(I)=k$, II is unmixed.
In Subsection  5.4, we state our final results and their corollaries, including another meaningful result obtained along the way: Hilbert’s Syzygy Theorem, which only requires the regularity of the polynomial ring in addition to our formalization.
Theorem 1.10 (Hilbert’s Syzygy).
For a field $k$, $\operatorname{gldim}(k[x_{1},\cdots,x_{n}])=n$.
Finally, in section  6, we describe implementation issues–particularly the universe issues, related works, and directions for future works.
@@
Throughout the paper, we use the symbol W for external links. Most statements and definitions are accompanied by such a link directly to the source code for the corresponding statement in mathlib or in a branch of our fork W. To keep the links stable, the links to the latter point to a fixed commit of the master branch (shortly after cleanup of the project); while other links point to the documentation of mathlib. In some code excerpts the dot notation is exhibited: for example, for I of type Ideal R, the call Ideal.depth I M can be shortened to I.depth M.
@@ Acknowledgements
We thank Professor Liang Xiao for his preliminary review and for organizing the collaborative work that made this project possible. We are grateful to Dr. Jiedong Jiang for his helpful suggestions in revising the first version into the second. We also thank Dr. Huanhuan Yu for producing the blueprint of the formalization project. We are also deeply grateful to all the contributors and reviewers who have been involved in the development and update process of both Lean and mathlib, whose efforts have made our work possible. Special thanks go to Professor Joël Riou for developing the derived category and the $\operatorname{Ext}$ functor, which enabled the homological methods. This work is supported in part by National Key R&D Program of China grant 2024YFA1014000.
@@ 2. Preliminaries
In this section, we review some preliminary results that will be used in later proofs. All code shown in this section are due to the authors, except for those indicated in Subsections  2.1 and  2.5.
@@ 2.1. Associated Primes
At the start of this project, we only had the definition associatedPrimes W and basic facts about associated primes in mathlibW. Thanks to the work of Jinzhao Pan, we now have the finiteness of associated primes for finitely generated modules over Noetherian rings in mathlibW.
@@
Combined with the fact that the complement of the non-zero-divisors of a module equals the union of its associated primesW, this lemma allows us to apply the prime avoidance lemma when looking for a regular element in an ideal.
In addition to existing results in mathlib, we formalized results on associated primes of localized modules.W
Lemma 2.1.
For $S$ a multiplicative set of a Noetherian ring $R$ and $M$ a finitely generated $R$-module, one has $\mathrm{Ass}(S^{-1}M)=\mathrm{Ass}(M)\cap\operatorname{Spec}(S^{-1}R)$.
Denoting $R^{\prime}$ and $M^{\prime}$ as the localizations of $R$ and $M$ at $S$ respectively, the main statement is:
@@
Since the induced map $\operatorname{Spec}(R^{\prime})\to\operatorname{Spec}(R)$ is injective, this gives a full characterization of $\mathrm{Ass}(S^{-1}M)$.
Furthermore, applying the above to the localization at $p$, a minimal prime over $\mathrm{Ann}(M)$, since associated primes of $M_{p}$ exist and any such associated prime must be the maximal ideal of $R_{p}$ (because $p$ is minimal), we obtain the following lemma.W
Lemma 2.2.
For $M$ a finitely generated module over Noetherian ring $R$, minimal primes over $\mathrm{Ann}(M)$ lie in $\mathrm{Ass}(M)$.
@@ 2.2. Krull Dimension of Module
Recall that the Krull dimension of a module is the Krull dimension of its support viewed as a preordered set. We implement it in Lean4 in the same way.W
def supportDim : WithBot ℕ∞ := krullDim (Module.support R M)
Some basic results are formalized hereW.
We prove a lemma that will be frequently used later.W
Lemma 2.3.
For a finitely generated $R$-module $M$ and an $M$-regular element $x$ contained in the Jacobson radical of $\mathrm{Ann}(M)$, then $\dim(M/xM)+1=\dim(M)$.
@@
Proof.
Note that any $M$-regular element is not contained in any minimal prime over $\mathrm{Ann}(M)$. The following two propositions suffices.
Proposition 2.4.
For a finitely generated $R$-module $M$, quotienting by an element in the Jacobson radical of $\mathrm{Ann}(M)$ decreases the dimension by at most one.
Proof.
The formalization of this proposition follows [10, Proposition 2.3.4] via [10, Lemma 2.3.3]W. ∎
Proposition 2.5.
For a finitely generated $R$-module $M$, quotienting by an element decreases the dimension by at least one.
By induction on the length of a regular sequence, we obtain the following corollary of Lemma  2.3.W
Corollary 2.6.
If $R$ is local, for a finitely generated $R$-module $M$ and an $M$-regular sequence $(r_{1},\cdots,r_{k})$, $\dim(M/(r_{1},\cdots,r_{k}))+k=\dim(M)$.
@@ 2.3. Dimension of Polynomial Rings
As we need to work with polynomials over Cohen–Macaulay and regular rings, some results about the heights of ideals in polynomial rings over Noetherian rings are required. Let $R$ be a Noetherian ring. We establish the following results, which follow easily from Krull’s height theoremW.
Proposition 2.7.
For local ring $(R,\mathfrak{m})$, $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})$
Proposition 2.8.
For ring $R$, let $\mathfrak{p}$ be a prime ideal of $R[X]$, then $0pt(\mathfrak{p})\leq 0pt(\mathfrak{p}\cap R)+1$
From these we obtain the standard result on the Krull dimension of polynomial rings over Noetherian rings.W
@@
There is an ongoing effort to formalize similar statements more generally by studying heights under ring homomorphisms with the going-down property.WW This would establish more general versions of the above ad hoc lemmas. These lemmas will be deprecated once the generalizations become available in mathlib. See Subsection  6.2 for details.
@@ 2.4. Regular Sequence Under Flat Base Change
Since we need to handle preservation of regular sequences under localization, we formalized the more general results that weakly regular sequences are preserved under flat base changeW and regular sequences are preserved under faithfully flat base changeW. These results also benefit the proof that polynomials over Cohen–Macaulay rings are again Cohen–Macaulay in Subsection  4.2.
@@
Specializing to the case where the module is the ring itself yields a more user-friendly version.WW
Finally, we also added specialized lemmas for localized modules and localization, which show that a regular sequence contained in the prime being localized at remains regular after localization.WW
@@ 2.5. Projective Dimension
Since the $\operatorname{Ext}$ functor was introduced to mathlib relatively recently as presented in [19], at the time of this work, besides the basic constructions inherited from shifted hom in derived category such as composition, we only had a predicate describing that an object $X$ of an abelian category $C$ has projective dimension strictly less than some $n\in\mathbb{N}$, defined as the vanishing of all $\operatorname{Ext}_{i}(X,Y)$ for all $Y$ and all $i\geq n$ (formalized by Professor Joël Riou).
@@
With only this definition, some operations are awkward (for example, adding a number to the projective dimension in the Auslander–Buchsbaum formula). Even if we try to pick out the minimal $n$ such that HasProjectiveDimensionLT X (n + 1), the disadvantages are significant: first, such a number may not exist at all (if $X$ does not have a finite projective resolution); second, finding it with Nat.find complicates operations and requires decidable instances. For these reasons, we want projective dimension as a concrete numeric object.
However, projective dimension cannot be a plain natural number: first, we need to give objects without finite projective resolution the value $\infty$; second, by common convention the projective dimension of the zero object is -$\infty$. We therefore define projective dimension in WithBot ℕ∞:
def projectiveDimension {C : Type u} [Category.{v, u} C] [Abelian C] (X : C) : WithBot ℕ∞ :=
@@
Although this definition looks somewhat verbose, it behaves according to the usual conventions. (Outside code excerpts we denote projectiveDimension by $\operatorname{projdim}$.)
With this definition, we can prove the equivalence projectiveDimension X < n HasProjectiveDimensionLT X n and many other useful lemmas to make the numeric version convenient in practice. Many lemmas related to HasProjectiveDimensionLT now have corresponding statements for projectiveDimension.
We also prove a lemma specific to the category of $R$-modules that is useful when reducing global dimension to the local case.W
@@
Proof.
For convenience in formalization, we did not directly prove that $\operatorname{Ext}$ commutes with localization. Noting that the left-hand side being -$\infty$ is equivalent to $M=0$ and similarly for the right-hand side -$\infty$, we instead prove that for each natural number $n$, the statement that the left-hand side is no greater than $n$ is equivalent to the right-hand side being no greater than $n$, by induction on $n$. We construct an exact sequence $0\rightarrow N\rightarrow F\rightarrow M\rightarrow 0$ with $F$ free, localizing at a prime $\mathfrak{p}$ yields an exact sequence $0\rightarrow N_{\mathfrak{p}}\rightarrow F_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}\rightarrow 0$. Using the existing lemma about projective dimension in short exact sequencesW, we reduce the case $n+1$ to the case $n$. We are then left to show that $M$ is projective if and only if each localization at prime $M_{\mathfrak{p}}$ is projective, which is already in mathlib WW. ∎
There is also a version of this lemma considering only maximal ideals. W
Similarly we prove the following lemma.W
@@
For $S$ a multiplicative set of $R$ and $R$-module $M$, projective dimension of $S^{-1}M$ over $S^{-1}R$ is no greater than projective dimension of $M$ over $R$.
lemma projectiveDimension_le_projectiveDimension_of_isLocalizedModule [Small.{v} R] (S : Submonoid R)
(M : ModuleCat.{v} R) : projectiveDimension (M.localizedModule S) projectiveDimension M